Two earthquake events happen exactly at the same time. Their locations are 100 km (Event-1) and
200 km (Event-1) away from your city, respectively. The wavelength of the transverse surface waves
generated from both the events is same and is λ = 432 m. While the frequencies of the Event-1 and
Event-2 are 10 and 38 Hz, respectively. (a) Which earthquake event is going to reach to you first?
(b) What would be the time difference between their arrival? [
1
Expert's answer
2020-05-28T13:14:53-0400

Find speed of waves:


"v=\lambda f,\\
v_1=432\cdot10=4320\text{ m/s},\\
v_2=432\cdot38=16416\text{ m/s}."

The time for the wave to reach the points:


"t_1=\frac{d_1}{v_1}=\frac{100000}{4320}=23\text{ s}.\\
\space\\
t_2=\frac{d_2}{v_2}=\frac{200000}{16416}=12\text{ s}."


The second event will occur first.

The time difference:


"\Delta t=t_1-t_2=23-12=11\text{ s}."